Hi There, I am trying to calculate the dividend payment in the formula below:-
18.95 = D/1.14 + D/1.14^2
It is for a homemade dividend question; and I am having trouble with the algebra.
Thank you kindly!

Answer by rothauserc(4718)

18.95 = D/1.14 + D/1.14^2
multiply both sides of = by 1.14^2
18.95*(1.14^2) = D*(1.14) + D
24.62742 = D*(1.14) + D
now factor D on right side of =
24.62742 = (1.14 +1)*D
divide both sides of = by 2.14
D = 11.51
now check result
18.95 = 11.51/1.14 + 11.51/1.14^2
18.95 = 10.10 + 8.86
18.95 approx equal 18.96
answer checks